# Multi-Level Embedding and Alignment Network with Consistency and Invariance Learning for Cross-View Geo-Localization

### v2

@@ IV-F Feature Distribution
To comprehensively evaluate the effectiveness of the MEAN model in cross-view geo-localization, we visualized the intra-class and inter-class distance distributions on the University-1652 dataset, as shown in Fig. 4(a-e). The analysis includes a comparison of two representative methods, MCCG and Sample4Geo, along with a baseline model to validate the contributions of each component in MEAN. Compared with the initial features (Fig. 4(a)), MCCG (Fig. 4(b)), Sample4Geo (Fig. 4(c)), and the baseline model (Fig. 4(d)), the MEAN model achieves substantial separation between intra-class and inter-class distances, resulting in more compact intra-class features and more distinctly separated inter-class features. However, both the baseline model and the MCCG exhibit limitations in intra-class compactness, inter-class separation, and cross-view consistency. Although Sample4Geo demonstrates improved discriminative ability and consistency over the baseline and MCCG, it still exhibits overlap in some areas, indicating suboptimal intra-class compactness and inter-class separation. By contrast, the MEAN model (Fig. 4(e)) significantly reduces intra-class distances while increasing inter-class distances, showcasing superior feature discriminability and cross-view consistency.
@@ IV-G Retrieval Results
In the satellite-to-drone task, the satellite perspective image serves as the query image, and MEAN efficiently matches the correct target in the drone perspective (green border). This outcome illustrates that MEAN maintains a high level of retrieval performance in cross-view geo-localization tasks, effectively capturing feature consistency across different viewpoints.
@@
V CONCLUSION
@@
In this paper, we propose a lightweight MAEN framework for cross-view geo-localization tasks. The framework aims to enhance feature representation capability and discriminability through a multi-branch structure, comprising progressive diversification embedding, global extension embedding, and cross-domain enhanced alignment. The progressive diversification embedding branch focuses on generating diverse feature embeddings to accommodate complex geographic view variations and utilizing contrastive learning to improve feature consistency and discriminability. The global extension embedding branch further optimizes the interaction between global and fine-grained features to realize the coherent expression of cross-domain information. The cross-domain enhanced alignment branch learns a shared mapping between domains through adaptive calibration and applies a cross-domain invariance alignment loss to overcome the limitations of relying on local detail alignment. This strengthens the intrinsic correlation of cross-domain features and further mines consistency and invariance within the embedding space. Experimental results demonstrate that MAEN achieves competitive performance to strike a notable balance between matching accuracy and computational efficiency, while exhibiting superior adaptability and robustness in cross-view geo-localization tasks and outperforming existing state-of-the-art methods in certain cases.
@@
In current cross-view geo-localization tasks, existing methods heavily rely on labeled paired images and label-driven supervised training. In the future work, we will explore a new self-supervised learning framework to alleviate the bottleneck of high data annotation costs in cross-view geo-localization.

### v4

@@ IV-F Feature Distribution
To comprehensively evaluate the effectiveness of the MEAN model in CVGL, we visualized the intra-class and inter-class distance distributions on the University-1652 dataset, as shown in Fig. 4 (a-e). The analysis includes a comparison of two representative methods, MCCG and Sample4Geo, along with a baseline model to validate the contributions of each component in MEAN. Compared with the initial features (Fig. 4(a)), MCCG (Fig. 4(b)), Sample4Geo (Fig. 4(c)), and the baseline model (Fig. 4(d)), the MEAN model achieves substantial separation between intra-class and inter-class distances, resulting in more compact intra-class features and more distinctly separated inter-class features. However, both the baseline model and the MCCG exhibit limitations in intra-class compactness, inter-class separation, and cross-view consistency. Although Sample4Geo demonstrates improved discriminative ability and consistency over the baseline and MCCG, it still exhibits overlap in some areas, indicating suboptimal intra-class compactness and inter-class separation. By contrast, the MEAN model (Fig. 4(e)) significantly reduces intra-class distances while increasing inter-class distances, showcasing superior feature discriminability and cross-view consistency.
@@ IV-G Retrieval Results
In the satellite-to-drone task, the satellite perspective image serves as the query image, and MEAN efficiently matches the correct target in the drone perspective (green border). This outcome illustrates that MEAN maintains a high level of retrieval performance in CVGL tasks, effectively capturing feature consistency across different viewpoints.
@@
V Discussion
The key difference between our proposed method and previous methods lies in the former’s ability to achieve a better trade-off between computational efficiency and model performance. CVGL is primarily deployed in resource-constrained scenarios like UAV platforms, where computational efficiency is of critical importance. State-of-the-art methods, such as CCR [44], Sample4Geo [5], SRLN [42], and DAC [22], often rely on high-performance feature extractors and complex module designs, and typically prioritize performance improvements at the expense of computational efficiency. As a result, they are less suitable for deployment in resource-constrained scenarios. Unlike existing methods that employ multi-head self-attention mechanisms in Transformer-based architectures [43] or deep convolutional networks [29] for feature extraction, we adopt a lightweight shallow backbone for initial feature extraction. We present a solution to address the inherent limitations of shallow architectures with suboptimal initial performance in cross-view feature representation. Through multiple embedding generation and fusion, global embedding representation learning, and cross-domain enhanced alignment, our method effectively captures cross-view structural and spatial representations without significantly increasing the parameter scale.
It should be noted that our method achieves excellent performance using fewer parameters. The lightweight 9-layer shallow backbone of a compact design significantly reduces the parameter scale and architectural complexity and achieves a 67% reduction in the parameter count compared with the methods [13, 5, 22, 11]. An efficient local and global feature enhancement strategy is proposed utilizing two parallel branches instead of the traditional deep modules and the attention mechanism to counteract the efficiency degradation induced by the shallow backbone. These two simple branches significantly enhance the semantic completeness and discrimination of shallow features with a lightweight network architecture. The effectiveness of the two branches benefits from the design of dilation rates in the dilated convolutions. We test different combinations of dilation rates and ultimately select a configuration that obtains the best receptive field coverage and semantic representation capability within the multi-scale structure. In addition, the parallel branches offer flexible optimization space to independently modulate the learning process, which promote more efficient parameter convergence toward the optimal solution of the loss functions. Furthermore, in the third branch, we incorporate a dimensionality reduction strategy at intermediate layers to lower computational expense and preserve the high- and low-dimensional spatial information. Ultimately, our method achieves competitive or superior performance with a lower computational overhead.
VI CONCLUSION
@@
In this paper, we propose a lightweight MAEN framework for CVGL tasks. The framework aims to enhance feature representation capability and discriminability through a multi-branch structure, comprising progressive diversification embedding, global extension embedding, and cross-domain enhanced alignment. The progressive diversification embedding branch focuses on generating diverse feature embeddings to accommodate complex geographic view variations and utilizing contrastive learning to improve feature consistency and discriminability. The global extension embedding branch further optimizes the interaction between global and fine-grained features to realize the coherent expression of cross-domain information. The cross-domain enhanced alignment branch learns a shared mapping between domains through adaptive calibration and applies a cross-domain invariance alignment loss to overcome the limitations of relying on local detail alignment. This strengthens the intrinsic correlation of cross-domain features and further mines consistency and invariance within the embedding space. Experimental results demonstrate that MAEN achieves competitive performance to strike a notable balance between matching accuracy and computational efficiency, while exhibiting superior adaptability and robustness in CVGL tasks and outperforming existing state-of-the-art methods in certain cases.
@@
In current CVGL tasks, existing methods heavily rely on labeled paired images and label-driven supervised training. In future work, we will explore a new self-supervised learning framework to alleviate the bottleneck of high data annotation costs in CVGL.